# MonoPCC: Photometric-invariant Cycle Constraint for Monocular Depth Estimation of Endoscopic Images

### v3

@@ Abstract
Photometric constraint is indispensable for self-supervised monocular depth estimation. It involves warping a source image onto a target view using estimated depth&pose, and then minimizing the difference between the warped and target images. However, the endoscopic built-in light causes significant brightness fluctuations, and thus makes the photometric constraint unreliable. Previous efforts only mitigate this relying on extra models to calibrate image brightness. In this paper, we propose MonoPCC to address the brightness inconsistency radically by reshaping the photometric constraint into a cycle form. Instead of only warping the source image, MonoPCC constructs a closed loop consisting of two opposite forward-backward warping paths: from target to source and then back to target. Thus, the target image finally receives an image cycle-warped from itself, which naturally makes the constraint invariant to brightness changes. Moreover, MonoPCC transplants the source image’s phase-frequency into the intermediate warped image to avoid structure lost, and also stabilizes the training via an exponential moving average (EMA) strategy to avoid frequent changes in the forward warping. The comprehensive and extensive experimental results on four endoscopic datasets demonstrate that our proposed MonoPCC shows a great robustness to the brightness inconsistency, and exceeds other state-of-the-arts by reducing the absolute relative error by at least 7.27%, 9.38%, 9.90% and 3.17%, respectively.
@@ 1 Introduction
    We conduct comprehensive and extensive experiments on four public endoscopic datasets, i.e., SCARED [1], SimCol3D [21], SERV-CT [5], and Hamlyn [17, 26, 20] and a public natural dataset, i.e., KITTI [7]. The comparison results with eight state-of-the-art methods demonstrate the superiority of MonoPCC by decreasing the absolute relative error by at least 7.27%, 9.38%, 9.90% and 3.17% on the four endoscopic datasets, respectively, and its strong ability to resist inconsistent brightness in the training. Moreover, the comparison results on KITTI further verify the competitiveness of MonoPCC even for the natural scenario, where the brightness changes are often not that significant.
[…]
@@ 3.4 EMA to Stabilize Forward Warping
3.5 Implementation Details
We train MonoPCC using a single NVIDIA RTX A6000 GPU. We utilize MonoViT and ResNet18 pre-trained on ImageNet[3] for DepthNet and PoseNet initialization, respectively. We set the batch size to 12, and use AdamW [14] as the optimizer. The learning rate decay is exponential and set to 0.9. The number of epochs and learning rate differ in the warm-up and follow-up training phases. In warm-up training, the number of epochs is set to 20 epochs, and the initial learning rate is set to $1\times 10^{-4}$ for DepthNet’s encoder and $5\times 10^{-5}$ for the rest weights. In follow-up training, the number of epochs is set to 10 epochs, and the initial learning rate is set to $5\times 10^{-5}$ for all network weights.
4 Experimental Settings
4.1 Comparison Methods
@@
Figure 5: The Abs Rel error maps of comparison methods on SCARED and SimCol3D. Opencv Jet Colormap is used for visualization.
@@ 4.2 Datasets
Since the original data is binocular, we only use the left view to simulate our focusing monocular situation. We split the dataset into 21,066 frames for training and 551 for test, which are consistent with the previous SOTA [24]. No videos are cross-used in both training and test.
@@ 5.1 Comparison with State-of-the-arts
Figure 6: The Abs Rel error maps of comparison methods on SERV-CT and Hamlyn. Opencv Jet Colormap is used for visualization.
@@ 5.1.1 Evaluation on SCARED and SimCol3D
Figure 7: The Abs Rel error maps of seven ablation variants, including effectiveness of three components. (a)-(g) correspond to the $1^{st}$ to the $7^{th}$ rows in Table 4. Opencv Jet Colormap is used for visualization.
@@ 5.3 Robustness to Severe Brightness Inconsistency
Figure 8: The Abs Rel error maps of MonoPCC and other similar modules against photometric inconsistency. (a)-(d) correspond to the $1^{st}$ to the $4^{th}$ rows in Table 5. Opencv Jet Colormap is used for visualization.
@@ 6 Conclusion
Self-supervised monocular depth estimation is challenging for endoscopic scenes due to the severe negative impact of brightness fluctuations on the photometric constraint. In this paper, we propose a cycle-form warping to naturally overcome the brightness inconsistency of endoscopic images, and develop a MonoPCC for robust monocular depth estimation by using a re-designed photometric-invariant cycle constraint. To make the cycle-form warping effective in the photometric constraint, MonoPCC is equipped with two enabling techniques, i.e., structure transplant module (STM) and exponential moving average (EMA) strategy. STM alleviates image detail degradation to validate the backward warping, which uses the result of forward warping as input. EMA bridges the learning of network weights in the forward and backward warping, and stabilizes the intermediate warped image to ensure an effective convergence. The comprehensive and extensive comparisons with $8$ state-of-the-arts on five public datasets, i.e., SCARED, SimCol3D, SERV-CT, Hamlyn, and KITTI, demonstrate that MonoPCC achieves a superior performance by decreasing the absolute relative error by at least 7.27%, 9.38%, 9.90% and 3.17% on the four endoscopic datasets, respectively, and shows the competitiveness even for the natural scenario. Additionally, two ablation studies are conducted to confirm the effectiveness of three developed modules and the advancement of MonoPCC over other similar techniques against brightness fluctuations. However, the current pipeline only leverages a single frame to infer the depth map, which is likely to lack context information and geometric consistency. In the future work, we plan to extend MonoPCC into the multi-frame based monocular depth estimation.

### v4

@@ Abstract
Photometric constraint is indispensable for self-supervised monocular depth estimation. It involves warping a source image onto a target view using estimated depth&pose, and then minimizing the difference between the warped and target images. However, the endoscopic built-in light causes significant brightness fluctuations, and thus makes the photometric constraint unreliable. Previous efforts only mitigate this relying on extra models to calibrate image brightness. In this paper, we propose MonoPCC to address the brightness inconsistency radically by reshaping the photometric constraint into a cycle form. Instead of only warping the source image, MonoPCC constructs a closed loop consisting of two opposite forward-backward warping paths: from target to source and then back to target. Thus, the target image finally receives an image cycle-warped from itself, which naturally makes the constraint invariant to brightness changes. Moreover, MonoPCC transplants the source image’s phase-frequency into the intermediate warped image to avoid structure lost, and also stabilizes the training via an exponential moving average (EMA) strategy to avoid frequent changes in the forward warping. The comprehensive and extensive experimental results on five datasets demonstrate that our proposed MonoPCC shows a great robustness to the brightness inconsistency, and exceeds other state-of-the-arts by reducing the absolute relative error by 7.27%, 9.38%, 9.90% and 3.17% on four endoscopic datasets, respectively; superior results on the outdoor dataset verify the competitiveness of MonoPCC for the natural scenario.
@@ 1 Introduction
    We conduct comprehensive and extensive experiments on four public endoscopic datasets, i.e., SCARED [1], SimCol3D [29], SERV-CT [5], and Hamlyn [25, 37, 28] and a public natural dataset, i.e., KITTI [8]. The comparison results with eight state-of-the-art methods demonstrate the superiority of MonoPCC by decreasing the absolute relative error by 7.27%, 9.38%, 9.90% and 3.17% on four endoscopic datasets, respectively, and its strong ability to resist inconsistent brightness in the training. Moreover, the comparison results on KITTI further verify the competitiveness of MonoPCC even for the natural scenario, where the brightness changes are often not that significant.
@@ 2.2 Brightness Inconsistency
2.3 Cycle Warping in Medical Image Analysis
Cycle warping is a crucial technique in the context of image processing, and its applications have been explored in various domains. The core concept behind cycle constraint is to apply the same operation twice in opposite directions, ensuring cycle consistency between input and output results. Early work on this concept was applied in dense correspondence tasks [53], where a 3D computer aided design (CAD) model was used to generate two synthetic views of paired images, and cycle consistency was enforced between these two synthetic views. Since then, cycle constraint has been extended to other tasks, such as video interpolation [31], object tracking [47], and label propagation [7]. More recently, cycle constraint has been employed in medical imaging scenarios with impressive results [17, 16, 22, 11]. For example, to preserve original topology during deformable registration, CycleMorph [17] trained two registration networks to generate forward and reverse directional deformation vector fields, yielding the cycle consistency between the reversed image and the original image. Gong et al. [11] introduced a cycle constraint-based approach for modeling surgical soft-tissue deformation. By measuring the similarity between warped pixels throughout the entire forward-backward cycle, this method demonstrated robust deformation modeling across various tissue patterns, making it particularly useful for surgical planning and simulation. However, directly applying cycle warping to our specific task would lead to significant side effects, such as loss of detail and increased training difficulty, due to the twice-warping operation.
2.4 Fast Fourier Transform in Medical Image Analysis
The Fast Fourier Transform (FFT) is a powerful technique that converts signals from the spatial domain to the frequency domain and has been widely applied in medical imaging tasks. For instance, FFT is commonly used in CT reconstruction [34] and MRI denoising [44]. Recent advances in deep learning have further extended the application of FFT to other areas, particularly for addressing domain shifts, such as domain generalization [20, 50], domain adaptation [45, 38], and style transfer [23]. In the context of domain generalization, FedDG [20] was developed to improve the generalizability of federated learning models across unseen domains. By extracting low-level distribution information in the frequency space of client samples and performing continuous interpolation between local and diverse domain samples, FedDG enables each local client to acquire information from multiple domains, enhancing the model’s robustness. To alleviate the spatial misalignment in atlas-based segmentation, Lv et al. [23] proposed an image-aligned style transformation to generate image-mask pairs with diverse patterns. Specifically, Fourier amplitude components between the target image and warped atlas were mixed, while the phase component of the warped atlas was retained. Inspired by Lv et al. [23], we align source-target image pairs before Structure Transplant Module (STM), thereby mitigating the risk of potential artifacts. Conversely, we ensure the preservation of the brightness pattern by retaining the amplitude components while swapping the phase components to reconstruct blurred details caused by warping.
@@ 3.4 EMA to Stabilize Forward Warping
3.5 Implementation Details
We train MonoPCC using a single NVIDIA RTX A6000 GPU. We utilize MonoViT and ResNet18 pre-trained on ImageNet[3] for DepthNet and PoseNet initialization, respectively. We set the batch size to 12, and use AdamW [21] as the optimizer. The learning rate decay is exponential and set to 0.9. The number of epochs and learning rate differ in the warm-up and follow-up training phases. In warm-up training, the number of epochs is set to 20 epochs, and the initial learning rate is set to $1\times 10^{-4}$ for DepthNet’s encoder and $5\times 10^{-5}$ for the rest weights. In follow-up training, the number of epochs is set to 10 epochs, and the initial learning rate is set to $5\times 10^{-5}$ for all network weights. The weights of estimation networks in the EMA-copy path are updated every 200 steps.
4 Experimental Settings
4.1 Comparison Methods
@@
Figure 5: The Abs Rel error maps of comparison methods on SCARED and SimCol3D, with close-up details highlighted. The regions of interest (ROIs) are outlined with red dashed lines, and the Opencv Jet Colormap is used for visualization.
@@ 4.2 Datasets
Since the original data is binocular, we only use the left view to simulate our focusing monocular situation. We split the dataset into 21,066 frames (from 28 videos) for training and 551 (from the rest 7 videos) for test, which are consistent with the previous SOTA [33]. No videos are cross-used in both training and test. Also, two trajectories from the 7 videos are used to evaluate the performance of pose estimation, including 410 and 833 frames, respectively.
@@ 4.3 Evaluation Metrics
To evaluate the performance of camera pose estimation, we follow a protocol similar to that used for depth estimation. Specifically, we align the predicted camera poses to the scale of the ground truth trajectory, and we adopt the 5-frame evaluation rule [52], using the absolute trajectory error (ATE) as the evaluation metric.
@@ 5.1 Comparison with State-of-the-arts
Figure 6: The Abs Rel error maps of comparison methods on SERV-CT and Hamlyn, with close-up details highlighted. The regions of interest (ROIs) are outlined with red dashed lines, and the Opencv Jet Colormap is used for visualization.
@@ 5.1.1 Evaluation on SCARED and SimCol3D
Figure 7: The Abs Rel error maps of seven ablation variants, including effectiveness of three components, with close-up details highlighted. (a)-(g) correspond to the $1^{st}$ to the $7^{th}$ rows in Table 4. The regions of interest (ROIs) are outlined with red dashed lines, and the Opencv Jet Colormap is used for visualization.
@@ 5.3 Robustness to Severe Brightness Inconsistency
Figure 8: The Abs Rel error maps of MonoPCC and other similar modules against photometric inconsistency, with close-up details highlighted. (a)-(d) correspond to the $1^{st}$ to the $4^{th}$ rows in Table 5. The regions of interest (ROIs) are outlined with red dashed lines, and the Opencv Jet Colormap is used for visualization.
@@ 5.4 Results on KITTI Benchmark
Table 7: Quantitative comparison results (Absolute Trajectory Error) of pose estimation on two trajectories of SCARED. The best results are marked in bold and the second-best underlined.
Trajectory.1 $\downarrow$ Trajectory.2 $\downarrow$ Average ATE $\downarrow$
Monodepth2 0.0748 0.0536 0.0606
FeatDepth 0.0767 0.0468 0.0567
HR-Depth 0.0747 0.0609 0.0654
DIFFNet 0.0760 0.0571 0.0633
Endo-SfMLearner 0.0754 0.0518 0.0596
AF-SfMLearner 0.0737 0.0464 0.0554
MonoViT 0.0760 0.0523 0.0601
Lite-Mono 0.0758 0.0477 0.0570
MonoPCC(Ours) 0.0730 0.0459 0.0548
Figure 11: The qualitative pose estimation comparison based on two SCARED trajectories.
5.5 Performance of Pose Estimation
In this subsection, we also compare the performance of ego-motion estimation between MonoPCC and other SOTA methods.
Table 7 presents the quantitative comparison of pose estimation on the two trajectories. While most methods, including MonoPCC, share the same PoseNet architecture, the performance of pose estimation varies significantly across methods. Notably, MonoPCC outperforms other SOTAs in pose estimation for both trajectories. Moreover, our PoseNet achieves an 8.82% relative reduction in ATE compared to the baseline MonoViT ($p<0.001$), a 1.08% relative reduction compared to AF-SfMLearner ($p=0.095$).
Fig. 11 shows the predicted trajectories of our method (MonoPCC) and five other SOTAs, for two trajectories of SCARED in 3D space. The trajectories are aligned at the starting point, and the positions of the endpoints indicate how well the trajectories match the ground truth. As can be seen, the endpoints of the MonoPCC trajectory (the last column) are closer to the ground truth (the blue curve), demonstrating the superiority of our approach in pose estimation.
Figure 12: Qualitative comparison results on the 3D scene reconstruction based on the estimated depth maps of two methods. The three sequences are selected from SCARED.
5.6 Performance of Scene Reconstruction
Using the known camera intrinsics and the estimated depth maps from continuous video frames, we aim to recover 3D surface models and merge them into a larger scene with richer spatial information. Specifically, we estimate the depth maps of three sequences from the SCARED dataset using the top-performing methods: MonoPCC and AF-SfMLearner. Following the setting of Endo-Depth-and-Motion [30], we apply photometric tracking and volumetric reconstruction to fuse the depth maps and generate 3D mesh models.
Fig. 12 presents three examples of the reconstructed scenes, along with close-up views of the details. In general, both depth estimation methods provide relatively good results for 3D reconstruction. However, by comparing the highlighted regions in Fig. 12, we observe that the surface meshes generated by AF-SfMLearner exhibit noticeable artifacts. These artifacts are likely due to depth outliers in the predicted maps. In contrast, our method (MonoPCC) produces smoother and more accurate 3D reconstructions, demonstrating its potential for high-quality 3D scene reconstruction.
@@ 6 Conclusion
Self-supervised monocular depth estimation is challenging for endoscopic scenes due to the severe negative impact of brightness fluctuations on the photometric constraint. In this paper, we propose a cycle-form warping to naturally overcome the brightness inconsistency of endoscopic images, and develop a MonoPCC for robust monocular depth estimation by using a re-designed photometric-invariant cycle constraint. To make the cycle-form warping effective in the photometric constraint, MonoPCC is equipped with two enabling techniques, i.e., structure transplant module (STM) and exponential moving average (EMA) strategy. STM alleviates image detail degradation to validate the backward warping, which uses the result of forward warping as input. EMA bridges the learning of network weights in the forward and backward warping, and stabilizes the intermediate warped image to ensure an effective convergence. The comprehensive and extensive comparisons with $8$ state-of-the-arts on five public datasets, i.e., SCARED, SimCol3D, SERV-CT, Hamlyn, and KITTI, demonstrate that MonoPCC achieves a superior performance by decreasing the absolute relative error by 7.27%, 9.38%, 9.90% and 3.17% on four endoscopic datasets, respectively, and shows the competitiveness even for the natural scenario. Additionally, two ablation studies are conducted to confirm the effectiveness of three developed modules and the advancement of MonoPCC over other similar techniques against brightness fluctuations.
Limitations. The current pipeline relies on a single frame to infer the depth map. Since each prediction is made independently, the model lacks perception of temporal consistency, meaning that the depth values at the same location may vary over time. This temporal inconsistency can lead to artifacts, such as overlapping tissue surfaces, as shown in the 3D reconstruction visualization in Fig. 12. Furthermore, our method is primarily designed for static endoscopic scenes. In dynamic scenarios involving tissue deformation, MonoPCC may not perform effectively. This is because the depth values at corresponding positions between source-target paired images can change locally, making it difficult to establish the cycle warping path consistently.
Potential Future Application. In this paper, we have demonstrated the effectiveness of the PCC strategy for self-supervised monocular depth estimation in endoscopic images. We believe that our framework can be seamlessly integrated into other related tasks, such as stereo matching [35] and metric depth estimation [42, 43], both of which face challenges due to brightness fluctuations. Additionally, in the field of NeRF-based scene reconstruction, several depth-prior-assisted methods [39, 19, 14] utilize estimated depth to guide model training. Therefore, our depth estimator, designed specifically for endoscopic scenes, could also enhance the performance of such downstream tasks.